# Closing the Data-Efficiency Gap Between Autoregressive and Masked Diffusion LLMs

### v1

@@ 3 Datasets and experimental setups
We investigate the data efficiency of dLLMs for knowledge injection by fine-tuning, focusing on whether they require paraphrases to successfully handle both forward and backward QAs. We follow the original pretraining protocol (Nie et al., 2025b) to fine-tune LLaDA-8B-Instruct on the dataset samples using the loss defined in Eq. 1. On three datasets, the accuracy difference between fine-tuning with and without paraphrases is much smaller in the dLLM than in the arLLM (Table 3): dLLM without paraphrases can already achieve decent accuracies on both forward and backward questions; fine-tuning with paraphrases can further increase the accuracy by a small amount. The accuracy difference between forward and backward questions is also smaller, indicating dLLM does not rely on paraphrases to answer backward questions. These results together suggest that dLLM has superior data efficiency and is free of the reversal curse in the post-training phase. By plotting the testing accuracy across the training steps (Figure 2), we observe that arLLM fine-tuned without paraphrases improves QA accuracy only in the beginning of training, then quickly decreases, indicating overfitting. The dLLM without paraphrases, on the other hand, does not show signs of overfitting. This finding echoes what has been found in comparing arLLMs and dLLMs in the pre-training phase (Prabhudesai et al., 2025; Ni & the team, 2025).
@@
One may expect that fine-tuning dLLM converges slower than arLLM, because learning any-order factorization requires seeing more than one way of the factorizations (i.e., samples masked in different ways) (Xue et al., 2025; Kim et al., 2025). However, we found that dLLM converges as fast as arLLM (Figure 2, Appendix Table 4); in the Biography dataset, dLLM even converges faster than arLLM. This indicates that dLLM does not trade better data efficiency and performance for more computations; it requires the same or less computation and fewer training samples, but achieves better downstream performance.
@@
Figure 2: Training dynamics of arLLM, dLLM, and masked arLLM. For the NameDescription dataset, forward and backward accuracy are the average of N2D and D2N types. Paraphrases used in the Wiki dataset are the same-order paraphrases set. Due to the randomness of sampling the masks, we average across 4 random seed for the dLLM and masked arLLM on NameDescription and Biography Datasets. Curves for each seed are shown in Appendix Figure 7 8.
@@ 6 Masked Fine-tuning of arLLM
Figure 3: An example of masked fine-tuning prompt. Random selection of text tokens are replaced by a [MASK] token. Tokens with yellow background are used to compute the autoregressive loss.
Inspired by the supremacy of dLLM in knowledge injection by fine-tuning, we attempt to adapt its advantages to arLLM. If an instruct arLLM is capable enough, one may prompt an arLLM to act like a dLLM. Specifically, given a masked document and an instruction to recover the original document, if the model has knowledge of the original document and the masked document contains sufficient cues to retrieve this knowledge, an instruct arLLM is supposed to respond with the correct original document. If the arLLM does not already have the knowledge of the original document, setting the ground truth document as the supervised fine-tuning target may implicitly teach the model that knowledge. We refer to this fine-tuning paradigm as “masked fine-tuning” of arLLM, and the resulting model as “masked arLLM”. Masked fine-tuning of arLLM, from a broader perspective, establishes a training objective similar to that of dLLMs, wherein the model learns to reconstruct the unmasked sequence from a masked input. Following the dLLM noise sampling strategy, we randomly replace sample tokens with a reserved special token during training, where the mask ratio $t$ is sampled from a uniform distribution $U(0.05,0.95)$. Note that each sample can be masked with a different mask ratio in different epochs. We evaluate the mask fine-tuned arLLM in the regular autoregressive way using the default chat template. The exact prompt used in the fine-tuning is provided in Figure 3, and more details can be found in Appendix A.4.
To demonstrate that the effectiveness of our masked fine-tuning is not due to a simple data augmentation effect, we conducted a control experiment that replaces the masked text in the prompt with random tokens (Appendix Figure 9). Using random tokens declines the accuracy of masked fine-tuning to a level comparable to that of naive arLLM fine-tuning. To confirm that the effectiveness of our masked fine-tuning stems from the objective itself and not merely a simple data augmentation effect (i.e., introducing varied input text prepended to the training sample), we conducted a control experiment. In this test, the masked document within the prompt was replaced entirely with random tokens (Appendix Figure 9). This substitution caused the accuracy of the masked fine-tuning to plummet to the level of naive arLLM fine-tuning.
@@ 7 Effects of fine-tuning mask ratio
Previous studies (Allen-Zhu & Li, 2024; 2025) claim that bidirectional BERT-like models struggle with even forward style knowledge extraction due to the mask loss, which causes the model to learn incorrect associations between tokens. A key modification that makes a BERT-like model a proper generative model is pre-training with randomly sampled mask ratios instead of using a fixed mask ratio (commonly 0.15 in BERT)(Nie et al., 2025b; Devlin et al., 2018). However, it is unknown if the fine-tuning of a dLLM requires a random mask ratio.
To investigate whether this necessity persists during post-training, we change the fine-tuning process of the dLLMs and masked arLLMs to use fixed mask ratios ($t$) instead of randomly sampling them during the training (Figure 4). Fine-tuning with some fixed mask ratios (0.75 and 0.5) can be as effective as the random mask ratio in knowledge injection. However, there is considerable performance variation across choices of $t$. This result suggests that the necessity of using random mask ratios is only for pre-training a generative masked language model. In the fine-tuning phase of this particular task domain, using a fixed mask ratio of around 0.75 is as effective as using random mask ratios.
@@ 8 Discussion
In this study, we find that dLLMs are more data-efficient for post-training knowledge injection than arLLMs, achieving high accuracy on both forward and backward style questions, even without paraphrase augmentation. In contrast, arLLMs depend heavily on paraphrases and struggle with backward questions, confirming the reversal curse. To bridge this gap, we introduce a masked fine-tuning paradigm for arLLMs that leverages diffusion-style mask reconstruction as an instruction tuning task without modifying the auto-regressive architecture or loss. The novel method allows arLLMs to reach near-perfect accuracy on forward and backward questions without relying on any paraphrases, closing the data efficiency gap between arLLMs and dLLMs. In summary, we provide an effective recipe to achieve new knowledge injection by fine-tuning in LLMs.
@@
We believe that such knowledge injection by fine-tuning will serve as a cornerstone for a self-evolving AI in the era of experience (Silver & Sutton, 2025). Engineering a dynamic memory system for LLMs is a trending research field, as agentic LLMs need to learn and evolve from their experiences (Zhang et al., 2025; Chhikara et al., 2025). Most current memory systems are based on external databases that store experiences and new knowledge as text. Such explicit textual memory has been successful due to the well-known in-context learning ability of LLMs. However, such memory systems have disadvantages as follows: 1) a limited context window and degradation of performance with long context Liu et al. (2023), 2) expensive computation due to long context, 3) difficulty in expressing implicit knowledge as text, such as knowledge of winning a chess game, and 4) the intrinsic limitation of using vector-based embeddings for retrieval (Weller et al., 2025). Parametric memory (i.e., memorizing by changing the network weights) does not have the above issues; however, due to the complications of fine-tuning an LLM, parametric memory is much less popular in production settings (Zhang et al., 2025). Furthermore, fine-tuning LLMs is shown to be inefficient in learning new factual knowledge (Ovadia et al., 2023; Mecklenburg et al., 2024; Gekhman et al., 2024; Soudani et al., 2024; Zhao et al., 2025; Lampinen et al., 2025), and has to rely on gating to overcome catastrophic forgetting Pan et al. (2025); Wang et al. (2024); Yu et al. (2023); Hartvigsen et al. (2023). Our study shows the feasibility of knowledge injection by fine-tuning through a mask recovery objective. These findings extend the known data efficiency of pre-training masked dLLMs (Prabhudesai et al., 2025; Ni & the team, 2025). The mask recovery objective uses a more flexible factorization, which can be seen as an implicit data augmentation. Therefore, it enables strong performance in both forward and backward style recalls without explicitly creating more paraphrases. Furthermore, we show that such fine-tuning data efficiency is not exclusive to dLLM and its bi-directional attention pattern, the same objective can be reformulated into a supervised fine-tuning task for arLLM. We show that training arLLMs with this novel paradigm closes the performance and data efficiency gap. This implies that one does not need to switch to a dLLM, but can use any of the existing arLLMs and still benefits from the data efficiency advantage.
A limitation of our study is its reliance on relatively simple datasets; real-life knowledge and incoming experiences are much more complex. Therefore, a promising future direction is to evaluate both dLLMs and our masked fine-tuning paradigm in more realistic agentic settings and on more complex data. We also see the future potential in extending our masked fine-tuning approach to other phases of LLM training, such as pre-training and reinforcement learning style training.
@@ A.2 LLM usage
We construct a Wiki from real Wikipedia articles following the protocol of Pan et al. (2025). We first crawl all the pages under the wiki category “Category:2025_by_month”, then filter out the pages that are created before January 1st, 2025. This process minimizes the leakage of this “new” knowledge to the base model. Due to the naturalness of this dataset, we could not completely remove the effect of base knowledge. Llada-Instruct has a slightly higher base model accuracy than Llama-3.1-8B-instruct, but they are qualitatively similar (Table 3). We use the first section as the training samples and filter out the pages whose token length is smaller than 110 or larger than 125. This results in 96 wiki articles. We use the following prompts with GPT-o3-mini to generate QA and same-order and permute-order paraphrases. We classify QAs into forward and backward styles. This is done by prompting GPT-o3-mini to generate keywords in the question and answer, then comparing their appearance order in the original text.
@@
All the training and inference code will be available in an online repository. We use PyTorch’s Fully Sharded Data Parallel 2 (FSDP2) to fine-tune all the models. We find that using mixed precision training is important for the fine-tuning performance (around 30% performance gain), and use the configs: MixedPrecisionPolicy(param_dtype="bf16", reduce_dtype="float32", cast_forward_inputs=True). All the experiments are full parameter fine-tuning on 4x 80G H100 GPUs. We use a batch size of 64 (16 per device) for all the experiments. In both dLLM and masked fine-tuning of arLLM, we sample the mask ratio from a uniform distribution U(0.05,0.95) for each batch (except for the fixed mask ratio experiments). Note that, unlike the original dLLM training recipes which use U(0,1) (Nie et al., 2025b), given that our sequence length is much shorter than the pre-training, we leave a small margin to avoid edge cases.
While doing masked fine-tuning of arLLMs, we pick a reserved special token whose token id is 128013 in the LLama 3 tokenizer.
@@
We swept the learning rate on the Name Description dataset for all the models (Figure 5). We choose to use learning rates that yield smooth gains in accuracy throughout the training process while achieving high final accuracy. The learning rate used in the main experiments is 5e-6 for arLLM; 1e-5 for dLLM; 3e-6 for masked arLLM.
@@
Figure 5: Learning rate sweep. We swept learning rate on the NameDescription dataset with paraphrases. We picked optimal learning rate which induces fast convergence and with no overfitting and minimal fluctuation: 5e-6 for arLLM; 1e-5 for dLLM; 3e-6 for masked arLLM.
Figure 6: Total accuracy (macro average of forward and backward accuracy). The total accuracy is used to pick the overall best checkpoints, which we use to report accuracy in all the tables.
Figure 7: Random seed effects in dLLM. Random seed determines the sampling of mask ratio and masked tokens. Each line represent a random seed.
Figure 8: Random seed effects in maksed arLLM. Random seed determines the sampling of mask ratio and masked tokens. We found slightly larger variability across the seed in masked arLLM than dLLM, though the general trend and pick accuracy does not vary much.
@@
Table 4: To compare the rate of convergence, we fit the accuracy curve as a function of training steps to $(1-e^{-kx}). “AA$" indicates accuracy at convergence; $k$ indicates rate of convergence.
A.5 Generation eamples

### v2

@@ 3 Datasets and experimental setups
We investigate the data efficiency of dLLMs for knowledge injection by fine-tuning, focusing on whether they require paraphrases to successfully handle both forward and backward QAs. We follow the original pretraining protocol (llada) to fine-tune LLaDA-8B-Instruct on the dataset samples using the loss defined in Eq. 1. On three datasets, the accuracy difference between fine-tuning with and without paraphrases is much smaller in the dLLM than in the arLLM (Table 1): dLLM without paraphrases can already achieve decent and comparable accuracies on both forward and backward questions; fine-tuning with paraphrases can further increase the accuracy by a small amount. These results together suggest that dLLM has superior data efficiency and is free of the reversal curse in the post-training phase. By plotting the testing accuracy across the training steps (Figure 2), we observe that arLLM fine-tuned without paraphrases improves QA accuracy only in the beginning of training, then quickly decreases, indicating overfitting. The dLLM without paraphrases, on the other hand, does not show signs of overfitting. This finding echoes what has been found in comparing arLLMs and dLLMs in the pre-training phase (prabhudesai2025diffusion; ni2025difflm).
@@
One may expect that fine-tuning dLLM converges slower than arLLM, because learning any-order factorization requires seeing more than one way of the factorizations (i.e., samples masked in different ways) (xue2025any; kim2025train). However, we found that dLLM converges at least as fast as arLLM (Figure 2, Table 4, Appendix Table 5); in the Biography dataset, dLLM even converges faster than arLLM. This indicates that dLLM does not trade better data efficiency and performance for more training compute; it requires the same or less training compute and fewer training samples but achieves better downstream performance.
@@
Figure 2: Training dynamics of arLLM (Llama 8B), dLLM (Llada), and masked arLLM (Llama 8B). For the NameDescription dataset, forward and backward accuracy are the average of N2D and D2N types. Paraphrases used in the Wiki dataset are the same-order paraphrases set. Due to the randomness of sampling the masks, we average across 4 random seed for the dLLM and masked arLLM on NameDescription and Biography Datasets. Curves for each seed are shown in Appendix Figure 10-11.
@@ 6 Masked Fine-tuning of arLLM
Figure 3: An example of masked fine-tuning prompt. Random selection of text tokens are replaced by a [MASK] token. Highlighted tokens are used to compute the autoregressive loss.
Inspired by the supremacy of dLLM in knowledge injection by fine-tuning, we attempt to adapt its advantages to arLLM. If an instruct arLLM is capable enough, one may prompt an arLLM to act like a dLLM. Specifically, given a masked document and an instruction to recover the original document, if the model has knowledge of the original document and the masked document contains sufficient cues to retrieve this knowledge, an instruct arLLM is supposed to respond with the correct original document. If the arLLM does not already have the knowledge of the original document, setting the ground truth document as the supervised fine-tuning target may implicitly teach the model that knowledge. We refer to this fine-tuning paradigm as “masked fine-tuning” of arLLM, and the resulting model as “masked arLLM”. Masked fine-tuning of arLLM, from a broader perspective, establishes a training objective similar to that of dLLMs, wherein the model learns to reconstruct the unmasked sequence from a masked input. Following the dLLM noise sampling strategy, we randomly replace sample tokens with a reserved special token during training, where the mask ratio $t$ is sampled from a uniform distribution $U(0.05,0.95)$. Note that each sample can be masked with a different mask ratio in different epochs. We evaluate the mask fine-tuned arLLM in the regular autoregressive way using the default chat template. The exact prompt used in the fine-tuning is provided in Figure 3 (more details in Appendix A.4), and the training objective formulation is:
$\mathcal{L}(\theta)=-\frac{1}{\sum m_{t}}\sum_{t=1}^{T}m_{t}\,\log p_{\theta}\!\big(s_{t}\,\big|\,s_{<t}\big)$ (2)
where $s$ is the constructed sequence which has the form of Figure 3; $m_{t}\in\{0,1\}$ selects which tokens are included in the loss which are 1 if the token is part of the original sample in the “assistant” window or it is an “end of sequence” token.
Overall, masked-finetuning of arLLM successfully inherits all the merits of the dLLM fine-tuning (Table 1, Figure 2). Masked arLLM surpasses arLLM fine-tuning in the pre-training style with a huge margin (Table 1) and achieves comparably high accuracy in both forward and backward question categories. Moreover, like dLLM, masked arLLM relies much less on paraphrases in the fine-tuning dataset to saturate the accuracy in most cases. The convergence rate of masked fine-tuning is also comparable to or in some cases faster than dLLMs (Figure 2, Table 4, Appendix Table 5), suggesting masked fine-tuning is data-efficient to achieve better down-stream QA tasks than traditional fine-tuning without costing additional training compute.
To confirm that the effectiveness of our masked fine-tuning stems from the objective itself and not merely a simple data augmentation effect (i.e., introducing varied input text prepended to the training sample), we conducted a control experiment, in which the masked document within the prompt was replaced entirely with random tokens (Appendix Figure 12). This substitution caused the accuracy of the masked fine-tuning to drop to the level of naive arLLM fine-tuning.
@@ 7 Effects of fine-tuning mask ratio
Previous studies (zhu24a; zhu24b) claim that bidirectional BERT-like models struggle with even forward style knowledge extraction due to the mask loss, which causes the model to learn incorrect associations between tokens. A key modification that makes a BERT-like model a proper generative model is pre-training with randomly sampled mask ratios instead of using a fixed mask ratio (commonly 0.15 in BERT)(llada; bert). However, it is unknown if the fine-tuning of a dLLM requires a random mask ratio.
To investigate whether this necessity persists during post-training, we change the fine-tuning process of the dLLMs and masked arLLMs to use fixed mask ratios ($t$) instead of randomly sampling them during the training (Figure 4). Fine-tuning with some fixed mask ratios (0.75 and 0.5) can be as effective as the random mask ratio in knowledge injection. However, there is considerable performance variation across the choices of $t$. Conceptually, varying $t$ modulates task difficulty; a moderate $t$ places training in a “hard-but-not-impossible” regime that maximizes the per-example gradient signal. The effectiveness of some fixed mask ratios indicate that a dLLM only needs to sample over t to learn the demasking process during pre-training. Once this ability is acquired and not forgotten during fine-tuning, a fixed $t$ is sufficient for learning new knowledge.
@@
8 Masked supervised fine-tuning
To study whether the advantages of masked fine-tuning over regular fine-tuning of arLLM extend beyond knowledge injection QA tasks, we test this new paradigm with SFT to improve the model’s math capability. Different from previous sections in which fine-tuning samples are raw text containing new knowledge, the SFT fine-tuning samples contain QA pairs.
Models GSM8K MATH
Llama 3b baseline 0.686 0.258
Llama 3b SFT 0.686 0.281
Llama 3b masked SFT (ours) 0.735 0.290
Qwen 4b baseline 0.591 0.174
Qwen 4b SFT 0.776 0.376
Qwen 4b masked SFT (ours) 0.789 0.379
Table 3: Standard versus masked SFT across two math datasets (evaluated with 0-shot, pass@1). See full model names in section 3.
Similarly to the prompt construction used in Section 6 and Figure 3, to leverage the demasking training objective in arLLMs, we first transform QA pairs into a demasking task. Specifically, the user prompt consists of the question, a randomly masked answer, and an instruction to recover the full answer; the assistant response is the original full answer (See Appendix 16 for an example of the prompt). The loss formulation is the same as Eq. 2, but the value of $m$ is different in masked SFT, which is $1$ when a token is in the assistant response and its corresponding token in the user prompt is masked, and 0 elsewhere. We test masked SFT on two models, i.e. Llama-3.2-3B-Instruct and Qwen3-4B-Instruct-250, and two popular math datasets, i.e. GSM8K and MATH. Each dataset contains their own training set and testing set. Under all the conditions, masked SFT surpass the traditional SFT (Table 3). Experimental details including learning rate and epoch sweep are in the Appendix 16.
9 Computational cost comparison across methods
We use the Wiki dataset to comprehensively characterize the computational cost of different training methods tested in Table 1 covering data preparation, training, inference and convergence to peak accuracy. Specifically, we test 8b parameter arLLM (Llama-3.1-8b-instruct) and dLLM (Llada) models in the following 5 conditions: (1) arLLM + pretraining style finetuning + w/o paraphrases, (2) arLLM + pretraining style finetuning + w paraphrases, (3) arLLM + pretraining style finetuning + reverse training (golovneva2024reversetrainingnursereversal), (4) dLLM + pretraining style finetuning + w/o paraphrases, (5) arLLM + masked SFT style finetuning + w/o paraphrases (ours) (Table 4).
Llama 8B Llada Masked Llama 8B
w/o paraphrases (1) w paraphrases (2) reverse training (3) w/o paraphrases (4) w/o paraphrases (5, ours)
Data paraphrase compute NA 0.1M tokens (GPT-o3-mini) NA NA NA
Training FLOPs theoretical $T$ $T$ $T$ $T$ $2T+c$
FLOPs empirical (TFLOPs/step) 91.9 92.7 91.9 101.6 199.3
wall time (s/step) 2.32 2.54 2.84 2.52 2.76
peak memory (GB) 37.9 37.9 37.9 37.9 51.1
Inference FLOPs theoretical $S$ $S$ $S$ $^{2} SS$
Convergence accuracy at convergence (Fwd) 0.241 0.630 0.344 0.897 0.933
accuracy at convergence (Bwd) 0.182 0.361 0.235 0.704 0.883
rate of convergence (Fwd) 0.0350 0.0069 0.0151 0.0052 0.0032
rate of convergence (Bwd) 0.1337 0.0130 0.0495 0.0081 0.0029
Table 4: Comparison of data preparation, training and inference computational costs among different model architecture and training methods on Wiki dataset. Bold indicates best single performance and underline indicates best tied performance.
Data Preparation
Only condition (2) requires computational expensive data preparation to generate semantically identical and natural paraphrases of the original dataset. For example, in the Wiki dataset, paraphrasing 10 per sample costs 0.1M generation tokens of GPT-o3-mini or equivalent models. This cost scales further with larger original datasets and the requirements for the diversity of paraphrases. In contrast, all other methods require no or very simple data transformations during the training (sampling mask or reverse sequence) without additional compute.
Training
We compare training FLOPs, wall time and peak allocated memory per step. Theoretical FLOPs calculation follows formulation introduced in kaplan2020scalinglawsneurallanguage. Specifically, the training cost of a dense transformer is approximated as $2N$ FLOPs per token for the forward pass and $4N$ for backpropagation, in total $6N$ FLOPs per token, where $N$ denotes the model parameter size. For a step that processes $T$ tokens in total, the theoretical per-step cost is therefore $\mathrm{FLOPs/step}\;\approx\;6\,N\,T$ In our comparisons, we keep the model size fixed at N=8N{=}8B and use the same effective global batch size across all conditions, so the only varying factor is the average sequence length of training samples. We thus report the theoretical quantity as a factor of $T$. Since paraphrasing or reverse training only modify the information order within the training samples, condition (1)-(4) have comparable average sequence length $T$. Condition (5) requires presenting the model with the original sequence and its masked counterpart in SFT format (Figure 3) doubling the average sequence length to $2T+c$ where $c$ is the constant accounting for SFT instructions. We also empirically measure the average FLOPs, wall time and peak memory per step using PyTorch Profiler (pytorch_profiler_docs). Empirical FLOPs measurements agree with the theoretical calculations. Among all conditions, our proposed method condition (5) costs around twice training FLOPs and slightly larger wall time as well as peak allocated memory. Additionally, we report that the sampling the mask and constructing the masked fine-tuning prompt during training for condition (4) and (5) take negligible time (<3.5ms) and memory comparing to the forward and backward pass, thus causing no significant overhead.
Inference
For inference statistics, we only compare the theoretical inference FLOPs as empirical values highly depend on the generation length of a particular answer. Following the same calculation detailed in the above section, an inference step produces a forward pass through the dense transformer, leading to $\mathrm{FLOPs/step}\;\approx\;2\,N\,S$ where $S$ denotes the average generation sequence length. Only condition (4) dLLM needs inference FLOPs quadratic in $S$ as generation in dLLM cannot reuse the KV cache since it changes after each denosing step (ni2025difflm). All other conditions are linear in $S$.
Convergence
[…]
@@ 10 Discussion
We believe that such knowledge injection by fine-tuning will serve as a cornerstone for a self-evolving AI in the era of experience (silver2025welcome). Engineering a dynamic memory system for LLMs is a trending research field, as agentic LLMs need to learn and evolve from their experiences (zhang2025survey; chhikara2025mem0). Most current memory systems are based on external databases that store experiences and new knowledge as text. Such explicit textual memory has been successful due to the well-known in-context learning ability of LLMs. However, such memory systems have disadvantages: 1) a limited context window and degradation of performance with long context liu2023lost, 2) expensive computation due to long context, 3) difficulty in expressing implicit knowledge as text, such as knowledge of winning a chess game, and 4) the intrinsic limitation of using vector-based embeddings for retrieval (weller2025theoretical). Parametric memory (i.e., memorizing by changing the network weights) does not have the above issues; however, due to the complications of fine-tuning an LLM, parametric memory is much less popular in production settings (zhang2025survey). Furthermore, fine-tuning LLMs is shown to be inefficient in learning new factual knowledge (ovadia2023fine; mecklenburg2024injecting; gekhman2024does; soudani2024fine; zhao2025style; lampinen2025generalization), and has to rely on gating to overcome catastrophic forgetting (pan2025memorization; wang2024wiserethinkingknowledgememory; yu2023meloenhancingmodelediting; hartvigsen2023aginggracelifelongmodel). Our study shows the feasibility of knowledge injection by fine-tuning through a mask recovery objective without additional paraphrases. These findings extend the known data efficiency of pre-training masked dLLMs (prabhudesai2025diffusion; ni2025difflm). The mask recovery objective uses a more flexible factorization, which can be seen as an implicit data augmentation. Therefore, it enables strong performance in both forward and backward style recalls. Furthermore, we show that such benefit is not exclusive to dLLM. The same objective can be reformulated into a supervised fine-tuning task for arLLM. We show that training arLLMs with this novel paradigm improves both knowledge injection and math SFT. This implies that one does not need to switch to a dLLM, but can use any of the existing arLLMs and still benefits from the data efficiency advantage.
@@ A.2 LLM usage
We construct a Wiki from real Wikipedia articles following the protocol of pan2025memorization. We first crawl all the pages under the wiki category “Category:2025_by_month”, then filter out the pages that are created before January 1st, 2025. This process minimizes the leakage of this “new” knowledge to the base model. Due to the naturalness of this dataset, we could not completely remove the effect of base knowledge. Llada-Instruct has a slightly higher base model accuracy than Llama-3.1-8B-instruct, but they are qualitatively similar (Table 1). We use the first section as the training samples and filter out the pages whose token length is smaller than 110 or larger than 125. This results in 96 wiki articles. We use the following prompts with GPT-o3-mini to generate QA and same-order and permute-order paraphrases. We classify QAs into forward and backward styles. This is done by prompting GPT-o3-mini to generate keywords in the question and answer, then comparing their appearance order in the original text.
@@
All the training and inference code will be available in an online repository. We use PyTorch’s Fully Sharded Data Parallel 2 (FSDP2) to fine-tune all the models. We find that using mixed precision training is important for the fine-tuning performance (around 30% performance gain), and use the configs: MixedPrecisionPolicy(param_dtype="bf16", reduce_dtype="float32", cast_forward_inputs=True). All the experiments are full parameter fine-tuning on 4x 80G H100 GPUs. We use a batch size of 64 (16 per device) for all the experiments. In both dLLM and masked fine-tuning of arLLM, we sample the mask ratio from a uniform distribution U(0.05,0.95) for each batch (except for the fixed mask ratio experiments). Note that, unlike the original dLLM training recipes which use U(0,1) (llada), given that our sequence length is much shorter than the pre-training, we leave a small margin to avoid edge cases.
While doing masked fine-tuning of Llama models, we pick a reserved special token whose token id is 128013 in the Llama tokenizer. While doing masked fine-tuning of Qwen models, we pick token “[]” whose token id is 1294 in the Qwen tokenizer.
@@
We use Adam optimizer with 0.1 weight decay coefficient; betas 0.9 and 0.95; 2% total steps as warm-up steps. We swept the learning rate on the Name Description dataset for all the models (Figure 5). We choose to use learning rates that yield smooth gains in accuracy throughout the training process while achieving high final accuracy. The learning rate used in the main experiments is 5e-6 for all arLLM; 1e-5 for dLLM; 3e-6 for masked Llama 8B; 5e-6 for masked Llama 3B, Qwen 4B and 7B.
@@
Figure 5: Learning rate sweep of Llama-3.1-8B-instruct. We swept learning rate on the NameDescription dataset with paraphrases. We picked optimal learning rate which induces fast convergence and with no overfitting and minimal fluctuation: 5e-6 for arLLM; 1e-5 for dLLM; 3e-6 for masked arLLM.
Figure 6: Total accuracy (macro average of forward and backward accuracy) of experiments on Llama-3.1-8B-Instruct. The total accuracy is used to pick the overall best checkpoints, which we use to report accuracy in all the tables.
Figure 7: Learning dynamics of Llama-3.2-3B-Instruct.
Figure 8: Learning dynamics of Qwen/Qwen3-4B-Instruct-2507.
Figure 9: Learning dynamics of Llama-3.2-3B-Instruct.
Figure 10: Random seed effects in Llada. Random seed determines the sampling of mask ratio and masked tokens. Each line represent a random seed.
Figure 11: Random seed effects in maksed Llama3.1 8B. Random seed determines the sampling of mask ratio and masked tokens. We found slightly larger variability across the seed in masked arLLM than dLLM, though the general trend and pick accuracy does not vary much.
@@
Table 5: To compare the rate of convergence, we fit the accuracy curve as a function of training steps to $(1-e^{-kx}). “AA$” is the accuracy at convergence; $k$ is the rate of convergence (unit $1/step$).
A.5 Masked SFT configs
In Section 8, we proposed a Masked SFT methods that turns any SFT prompt-response sequence into a demasking task. Specifically, the constructed demasking task sequence contains the question and and masked answer in the user prompt, and full answer in the assistant response. For example, an data entry in th GSM8K dataset is: Question: Joy can read 8 pages of a book in 20 minutes. How many hours will it take her to read 120 pages? Answer: In one hour, there are 3 sets of 20 minutes. So, Joy can read 8 x 3 = <<8*3=24>>24 pages in an hour. It will take her 120/24 = <<120/24=5>>5 hours to read 120 pages. 5. After generating a random mask, the constructed masked SFT sequence is:
[…]
The highlighted tokens are where the loss is calculated. More specifically, in Eq. 2, $m_{i}$ is set to $1$ when the $i$th token is in the assistant response and its corresponding token in the user prompt is masked.
We choose to use the GSM8K and MATH datasets for testing the SFT performance on Llama-3.2-3B-Instruct and Qwen/Qwen3-4B-Instruct-2507. For the MATH dataset, we use the default subset from huggingface DigitalLearningGmbH/MATH-lighteval. We further filter out the training samples whose token lengths (question + answer) are longer than 512. When evaluating the resulting models, we use the evaluation framework LM Evaluation Harness and the default tasks gsm8k and hendrycks_math (eval-harness). Specifically, we choose to use 0-shot and pass@1 with a maximum generation length of 256 at a temperature of 0. For GSM8K we report the accuracy using exact match with LM Evaluation Harness’s flexible extraction. For MATH we report the accuracy using exact match with math-verify extraction (Kydlicek_Math-Verify_Math_Verification). Both extraction methods are chosen to maximize alignment with human examination.
Most of the training configurations are the same as those in the main experiments, except for the following changes. The batch size is 32 for the GSM8K dataset, and 16 for the MATH dataset. We set maximum training epoch to 7, and reported the best testing accuracy across the training. Optimal learning rate is found for each model and dataset pair and shown in the following figures (Appendix Figure 13-16).
Figure 13: Learning rate and epoch sweep of Llama-3.2-3B-Instruct on GSM8K dataset.
Figure 14: Learning rate and epoch sweep of Qwen3-4B-Instruct-2507 on GSM8K dataset.
Figure 15: Learning rate and epoch sweep of Llama-3.2-3B-Instruct on MATH dataset.
Figure 16: Learning rate and epoch sweep of Qwen3-4B-Instruct-2507 on MATH dataset.
A.6 On reversal curse
Prior studies have justified the reversal curse as an intrinsic limitation of arLLM training (zhu2024towards; kitouni2024factorization; zhu2024towards). Here we provide an explanation that is conceptually easy to grasp. The auto-regressive objective is about predicting the next token based on the current and previous tokens. If the prediction of one next token requires a piece of new knowledge (i.e., it cannot be predicted based on the current knowledge in the weights or previous tokens), the loss will force the weights to change to favor such a prediction. More specifically, the change of weights induces a different representation (i.e., intermediate layer activations) of the previous tokens that favors the prediction of the next token. Since feedforward layers can be considered associative memory (meng2022locating), the change, conceptually, could be associating a new attribute with the representation of a token. However, such change does not affect the representation of future tokens to favor the prediction of the current token, since they do not contribute to the prediction of the “next” token. Thus, the future tokens could not learn a new association to its preceeding tokens. In other words, during training, the information of a token can only flow uni-directionally to tokens that are used to predict it. This has been named the “factorization curse" (kitouni2024factorization). This intuition can explain why the masked fine-tuning of arLLM resolves the curse. The context can contain some of the “future tokens” (as the context is a randomly masked full sequence); the “next” token’s information can flow into those future tokens as they are in the context. Paraphrases mitigates the reversal curse by creating views where the “future” cue becomes left context. For a fact $(x\!\to\!y)$ usually realized as “xy…x…y…”, paraphrases are fluent variants that put y before x (passive voice, topicalization, apposition, clause reordering), producing complementary left-to-right gradients and corpus-level bidirectional coverage without changing the AR objective. However, globally reversing all the information units in the documents can breaks discourse, coreference, and grammar, creating distribution shift.
A.7 Generation eamples